OpenStudy (anonymous):

please help me with this, I don't understand how to do this How can the properties of rational exponents be applied to simplify expressions with radicals or rational exponents?

OpenStudy (jdoe0001):

got expression to simplify?

OpenStudy (anonymous):

no i don't i think its just how to do it but if you have a example to show how to that would help aswell

OpenStudy (jdoe0001):

well... ok... one sec

OpenStudy (anonymous):

okay Thanks

OpenStudy (jdoe0001):

\(\large { \cfrac{\sqrt[3]{3^2}}{\sqrt[3]{3^4}}\impliedby \textit{expression with radicals} \\ \quad \\ \cfrac{\sqrt[3]{3^2}}{\sqrt[3]{3^4}}\implies \cfrac{3^{\frac{2}{3}}}{3^{\frac{4}{3}}}\implies \cfrac{3^{\frac{2}{3}}}{1}\cdot \cfrac{1}{3^{\frac{4}{3}}}\implies 3^{\frac{2}{3}}\cdot 3^{{\color{red}{ -}}\frac{4}{3}} \\ \quad \\ 3^{\frac{2}{3}-\frac{4}{3}}\implies 3^{-\frac{2}{3}}\implies \cfrac{1}{3^{\frac{2}{3}}}\implies \cfrac{1}{\sqrt[3]{3^2}} }\)